# thanks: These authors equally contributed to this workthanks: These authors equally contributed to this workthanks: A professor emeritus

### v2

@@ I Introduction
Recently, the generation of non-trivial mixed states having no counterpart of pure states attracts lots of interest in condensed matter physics as well as quantum information communities. As an example, a topologically-ordered pure state [8, 9] tends to change to a mixed state with another type of topological order [10, 11, 12, 13, 14, 15, 16]. From this point of view, behavior of symmetry protected topological (SPT) states under decoherence has been studied [17, 18, 19] to find that the SPT order survives in an ensemble level [20, 21]. In order to investigate these phenomena, we note that there are two types of symmetries; strong and weak symmetries in mixed states [22]. The notion of these symmetries can lead to some classification of mixed state orders. Then for mixed states, we have to reexamine notion of spontaneous symmetry breaking (SSB), that is, there are several types of SSBs in various systems, such as strong symmetry SSB, weak symmetry SSB, and strong to weak SSB (SWSSB), etc, [23, 24, 25, 26, 11, 12, 27, 28, 29, 30, 31, 32, 33] some of which are to be carefully defined in this work. Getting deep understanding of relation between various SSBs and discovering and proposing concrete examples of SSB phenomena induced by decoherence are currently one of the most important problems. In general, exact theoretical treatment of mixed states is not easy, and some of previous studies employed Choi isomorphism technique and the doubled Hilbert state formalism [34, 35]. By using these techniques as well as effective field theory methods, symmetry properties of certain specific mixed states are discussed [23, 17].
@@ III Doubled Hilbert space formalism
where, $\hat{I}_{j,u(\ell)}$ is an identity operator for site-$j$ vector space in $\mathcal{H}_{u(\ell)}$, $Z(X)_{j,u(\ell)}$ is Pauli-$Z$($X$) operator at site $j$ and $\tau_{zz(x)}=\tanh^{-1}[{p_{zz(x)}/(1-p_{zz(x)})}]$. Here, we note that the channel operator $\hat{\mathcal{E}}$ is not a unitary map in general cases although the channel is a CPTP map [23, 17]. Thus, the application of the channel operator generally changes the norm of the state vector.
@@ V Analysis of decohered state vectors in MPS formalism
For the practical numerical calculation, we search for the condition on the probabilities $(p_{zz},p_{x})$ to realize the decoherence corresponding to $H_{qAT}$ with $\lambda_{zz}=\lambda_{x}(=\lambda)$ for various values of $J/h$. From the parameter correspondence discussed in the previous section, we expect $J\lambda_{zz}=c\tau_{zz}(p_{zz})$ and $h\lambda_{x}=c\tau_{x}(p_{x})$, where $c$ is a positive constant and we set $h=1$, hereafter. After some algebra, we find that the conditions $\lambda_{zz}=\lambda_{x}$ and $(1/J)\tau_{zz}(p_{zz})=\tau_{x}(p_{x})$ are satisfied with $p_{x}=1/2-(1/2)(1-2p_{zz})^{1/J}$, which is a decreasing function of $p_{zz}$ for $J>0$. In the practical protocol, we vary the value of $p_{zz}$ and fix the corresponding value of $p_{x}$ using the above equation, and then apply the channel operations ^$\hat{\mathcal{E}}_{ZZ}(p_{zz})$ and ^$\hat{\mathcal{E}}_{X}(p_{x})$. It is obvious that this procedure preserves the condition $\lambda_{zz}=\lambda_{x}$ in the corresponding qAT model. Increase of $p_{zz}$ with the relation $p_{x}=1/2-(1/2)(1-2p_{zz})^{1/J}$ corresponds to an increase of $\lambda$ in the qAT model.
@@
Figure 4: Entanglement entropy for $J/h=0.8$ [(a)], $J/h=1$ [(b)] and $J/h=1.2$ cases [(c)]. Schematic entanglement cut for $L=8$ system under periodic boundary conditions is shown in the upper-left panel, where the orange dotted lines represent the entanglement cuts. Each insets show the extrapolation of data to estimate the transition probability of $p_{zz}$, denoted by $p_{zz}^{c}$, at which $S_{A}$ becomes maximum for $L\to\infty$. Here, to estimate locations of peaks of $S_{A}$, we employed a sixth-order polynomial function to estimate $p_{zz}^{c}$ for each system size. Using the data of finite size systems and a first-order polynomial function, $a/L+p^{c}_{zz}$, where $a$ and $p^{c}_{zz}$ are fitting parameters, we estimated the phase transition point in the thermodynamic limit.
@@ VI Numerical results by using MPS
We investigate the decohered (filtered) state in Eq. (1) as MPS by efficiently employing TeNPy package [50, 51]. In Eq. (1), the initial state $|\rho_{0}\rangle\rangle$, which is the ground state of the decoupled 1D TFIM in Fig. 1(b), is prepared by the DMRG. The filtering operation in Eq. (1) can be also efficiently carried out by the TeNPy package. We then obtain the state $|\rho_{D}\rangle\rangle$ for each probability parameters, $p_{zz}$ and $p_{x}$. The code reliability employed in this work is examined in Appendix C.
@@
The numerical results of $S_{A}$ are shown in Figs. 4(a)-4(c). For the data (a) ($J/h=0.8$ case), we find that data of $S_{A}$ exhibit a peak around $p_{zz}\sim 0.36$, and the system-size dependence of $S_{A}$ develops there, indicating the existence of a phase transition. For the locations of the observed peaks, we perform the linear extrapolation with respect to $1/L$ [See the inset in Fig. 4(a)], and obtain an estimation of a transition point as $p^{c}_{zz}\sim 0.372$ for $L\to\infty$. Similarly for both the other data (b) and (c) ($J/h=1$ and $1.2$ case), $S_{A}$ has a peak and the value of $S_{A}$ at the peak increases as the system size is getting larger. Thus, we think that the two cases also exhibit a phase transition around $p_{zz}\sim 0.30$ for the case (b) and $p_{zz}\sim 0.39$ for the case (c), respectively. As the case (a), by using the linear extrapolation [See the insets in Figs. 4(b) and  4(c)], we estimate the transition point for $L\to\infty$ as $p^{c}_{zz}\sim 0.306$ for $J/h=1$ case and $p^{c}_{zz}\sim 0.393$ for $J/h=1.2$ case, respectively. The existence of the above phase transitions is good agreement with the phase diagram of the qAT model. As an additional interesting numerical result, we display the subsystem-size dependence of the EE in Appendix D. In particular, estimation of the central charge of a possible CFT is given there.
@@ VII Summary of mixed state order from numerical calculation of correlators
Before going into summery of pattern of the SSB, we have to carefully examine how the $Z_{2}$ symmetry is realized in the initial state $\rho_{0}$. For $J/h>1$, the initial state $\rho_{0}$ obviously has a long-range order for the $Z_{2}$ parity. By setting $\rho_{0}$ to the $Z_{2}$ parity $\prod X_{j}=+1$ cat state, the initial state $\rho_{0}$ can be regarded as a strong symmetric state. In the numerical calculation, we employed this prescription as the system is large but still finite. On the other hand for $J/h<1$, the initial state $\rho_{0}$ is trivially strong symmetric under the $Z_{2}$ parity.
@@
As shown in the table I, the regime I has no specific character called trivial paramagnetic mixed state. The region II has non-trivial properties indicated by the numerical observation, i.e., it corresponds to the $Z_{2}$ SWSSB phase in the mixed state picture, and in the doubled Hilbert space picture, that is the state with the diagonal $Z_{2}$ symmetry as observed by the correlator $C^{u}_{ZZ}$. The region III is also non-trivial, i.e., shown by the numerical observation, that regime corresponds to the strong-to-trivial $Z_{2}$ SSB phase in the mixed state picture since both strong and weak SSB order parameters are finite. In the doubled Hilbert space picture, the state in the regime III exhibits $Z_{2}\times Z_{2}$ SSB, as the upper and lower chains have independent long-range order, which can be observed by the behavior of $C^{u}_{ZZ}$.
@@ VIII Summary and conclusion
In this work, utility and specific character of the multiple decoherence have been clarified. We expect that similar phenomena observed for the quantum Ising model will appear in other models under multiple decoherence, such as a 1D $Z_{2}$ gauge-Higgs model that attracts interest in the high-energy and condensed matter communities these days.
@@ Appendix B Canonical $Z_{2}$ correlator in the doubled Hilbert space
For the 1D TFIM under decoherence, the canonical $Z_{2}$ correlator is given by ${\rm Tr}[\rho_{D}Z_{i}Z_{j}]$, employed as an order parameter characterizing the ordinary long-range order in statistical mechanics. As we explained in the main text, this observable detects SSB of both the strong and weak symmetries. It is useful to have an expression corresponding to this in the doubled Hilbert space formalism to analyze mixed states. By using the Choi isomorphism formula for density matrix $\rho$  [34, 23],

### v3

@@ I Introduction
Recently the generation of non-trivial mixed states having no counterpart of pure states attracts lots of interest in condensed matter physics as well as quantum information communities. As an example, a topologically-ordered pure state [8, 9] tends to change to a mixed state with another type of topological order [10, 11, 12, 13, 14, 15, 16]. From this point of view, behavior of symmetry protected topological (SPT) states under decoherence has been studied [17, 18, 19] to find that the SPT order survives in an ensemble level [20, 21]. In order to investigate these phenomena, we note that there are two types of symmetries; strong and weak symmetries in mixed states [22]. The notion of these symmetries can lead to some classification of mixed state orders. Then for mixed states, we have to reexamine notion of spontaneous symmetry breaking (SSB), that is, there are several types of SSBs in various systems, such as strong symmetry SSB, weak symmetry SSB, and strong to weak SSB (SWSSB), etc, [23, 24, 25, 26, 11, 12, 27, 28, 29, 30, 31, 32, 33] some of which are to be carefully defined in this work. Getting deep understanding of relation between various SSBs and discovering and proposing concrete examples of SSB phenomena induced by decoherence are currently one of the most important problems. In general, exact theoretical treatment of mixed states is not easy, and some of previous studies employed Choi isomorphism technique and the doubled Hilbert state formalism [34, 35]. By using these techniques as well as effective field theory methods, symmetry properties of certain specific mixed states are discussed [23, 17].
@@ III Doubled Hilbert space formalism
where, $\hat{I}_{j,u(\ell)}$ is an identity operator for site-$j$ vector space in $\mathcal{H}_{u(\ell)}$, $Z(X)_{j,u(\ell)}$ is Pauli-$Z$($X$) operator at site $j$ and $\tau_{zz(x)}=\tanh^{-1}[{p_{zz(x)}/(1-p_{zz(x)})}]$. Here, we note that the channel operator $\hat{\mathcal{E}}$ is not a unitary map in general cases although the channel is a completely-positive trace-preserving map [40, 23, 17]. Thus, the application of the channel operator generally changes the norm of the state vector.
@@ V Analysis of decohered state vectors in MPS formalism
For the practical numerical calculation, we search for the condition on the probabilities $(p_{zz},p_{x})$ to realize the decoherence corresponding to $H_{qAT}$ with $\lambda_{zz}=\lambda_{x}(=\lambda)$ for various values of $J/h$. From the parameter correspondence discussed in the previous section, we expect $J\lambda_{zz}=c\tau_{zz}(p_{zz})$ and $h\lambda_{x}=c\tau_{x}(p_{x})$, where $c$ is a positive constant and we set $h=1$, hereafter. After some algebra, we find that the conditions $\lambda_{zz}=\lambda_{x}$ and $(1/J)\tau_{zz}(p_{zz})=\tau_{x}(p_{x})$ are satisfied with $p_{x}=1/2-(1/2)(1-2p_{zz})^{1/J}$, which is an increasing function of $p_{zz}$ for $J>0$. In the practical protocol, we vary the value of $p_{zz}$ and fix the corresponding value of $p_{x}$ using the above equation, and then apply the channel operations ^$\hat{\mathcal{E}}_{ZZ}(p_{zz})$ and ^$\hat{\mathcal{E}}_{X}(p_{x})$. It is obvious that this procedure preserves the condition $\lambda_{zz}=\lambda_{x}$ in the corresponding qAT model. Increase of $p_{zz}$ with the relation $p_{x}=1/2-(1/2)(1-2p_{zz})^{1/J}$ corresponds to an increase of $\lambda$ in the qAT model.
@@
Figure 4: Entanglement entropy for $J/h=0.8$ [(a)], $J/h=1$ [(b)] and $J/h=1.2$ cases [(c)]. Only for the case (b), we carried out the calculation up to $L=32$ system size for numerical accuracy since we start from the initial critical state. Schematic entanglement cut for $L=8$ system under periodic boundary conditions is shown in the upper-left panel, where the orange dotted lines represent the entanglement cuts. Panels in the bottom line show the extrapolation of data to estimate the transition probability of $p_{zz}$, denoted by $p_{zz}^{c}$, at which $S_{A}$ becomes maximum for $L\to\infty$. Here, to estimate locations of peaks of $S_{A}$, we employed a sixth-order polynomial function to estimate $p_{zz}^{c}$ for each system size. Using the data of finite size systems and a first-order polynomial function, $a/L+p^{c}_{zz}$, where $a$ and $p^{c}_{zz}$ are fitting parameters, we estimated the phase transition point in the thermodynamic limit.
@@
A concrete example of $L=8$ ladder system is shown in the upper-left panel in Fig. 4.
@@ VI Numerical results by using MPS
We investigate the decohered (filtered) state in Eq. (1) as MPS by efficiently employing TeNPy package [50, 51]. In Eq. (1), the initial state $|\rho_{0}\rangle\rangle$, which is the ground state of the decoupled 1D TFIM in Fig. 1(b), is prepared by the DMRG. In all numerical simulations in this work, we set maximum bond dimension $D=200$ and truncate the singular value less than $\mathcal{O}(10^{-6})$, and the energy convergence of the iterative DMRG sweeping is $\Delta E<\mathcal{O}(10^{-4})$ to obtain the initial MPS ground state. The filtering operation in Eq. (1) can be also efficiently carried out by the TeNPy package. We then obtain the state $|\rho_{D}\rangle\rangle$ for each probability parameters, $p_{zz}$ and $p_{x}$. The code reliability employed in this work is examined in Appendix C.
@@
The numerical results of $S_{A}$ are shown in Figs. 4(a)-4(c). For the data (a) ($J/h=0.8$ case), we find that data of $S_{A}$ exhibit a peak around $p_{zz}\sim 0.36$ 333 In the data in Fig. 4(a), around $p_{zz}\sim 0.1$, there is a dip of $S_{A}$, where there are no system-size dependence nor any signals of a phase transition. The reason for the dip of EE seems to be very complicated, and we think that it is induced by non-perturbative interplay between X+ZZ decoherences , and the system-size dependence of $S_{A}$ develops there, indicating the existence of a phase transition. For the locations of the observed peaks, we perform the linear extrapolation with respect to $1/L$ [See the bottom panel in Fig. 4(a)], and obtain an estimation of a transition point as $p^{c}_{zz}\sim 0.372$ for $L\to\infty$. Similarly for both the other data (b) and (c) ($J/h=1$ and $1.2$ case), $S_{A}$ has a peak and the value of $S_{A}$ at the peak increases as the system size is getting larger. Thus, we think that the two cases also exhibit a phase transition around $p_{zz}\sim 0.30$ for the case (b) and $p_{zz}\sim 0.39$ for the case (c), respectively. As the case (a), by using the linear extrapolation [See the bottom panels in Figs. 4(b) and  4(c)], we estimate the transition point for $L\to\infty$ as $p^{c}_{zz}\sim 0.308$ for $J/h=1$ case and $p^{c}_{zz}\sim 0.393$ for $J/h=1.2$ case, respectively. The existence of the above phase transitions is good agreement with the phase diagram of the qAT model. As an additional interesting numerical result, we display the subsystem-size dependence of the EE in Appendix D. In particular, estimation of the central charge of a possible CFT is given there.
@@ VII Summary of mixed state order from numerical calculation of correlators
Before going into summery of pattern of the SSB, we have to carefully examine how the $Z_{2}$ symmetry is realized in the initial state $\rho_{0}$. For $J/h>1$, the initial state $\rho_{0}$ obviously has a LRO for the $Z_{2}$ parity. By setting $\rho_{0}$ to the $Z_{2}$ parity $\prod X_{j}=+1$ cat state, the initial state $\rho_{0}$ can be regarded as a strong symmetric state. In the numerical calculation, we employed this prescription as the system is large but still finite. On the other hand for $J/h<1$, the initial state $\rho_{0}$ is trivially strong symmetric under the $Z_{2}$ parity.
@@
As shown in the table I, the regime I has no specific character called trivial paramagnetic mixed state. The region II has non-trivial properties indicated by the numerical observation, i.e., it corresponds to the $Z_{2}$ SWSSB phase in the mixed state picture, and in the doubled Hilbert space picture, that is the state with the diagonal $Z_{2}$ symmetry as observed by the correlator $C^{u}_{ZZ}$. The region III is also non-trivial, i.e., shown by the numerical observation, that regime corresponds to the strong-to-trivial $Z_{2}$ SSB phase in the mixed state picture since both strong and weak SSB order parameters are finite. In the doubled Hilbert space picture, the state in the regime III exhibits $Z_{2}\times Z_{2}$ SSB, as the upper and lower chains have independent LRO, which can be observed by the behavior of $C^{u}_{ZZ}$.
It is quite interesting and also instructive to find that the ground-state phase diagram of the qAT model [39] and that of the above mixed state emergent due to $X$ and $ZZ$ decoherence are quite similar. In fact, the regions I and III correspond to the paramagnetic state without LROs and the SSB state with the LRO, respectively. The properties of the above two regions and phases stem from the original TFIM, and therefore, the above similarity is somewhat natural. More interesting observation is that the region II of the SWSSB can be regarded as a counterpart of the spin-glass type phase II, which is generated by the sufficiently large $\lambda$-terms in $H_{qAT}$. In the qAT model, the $X$ and $ZZ$ operators in the $\lambda$-terms compete with each other, and as a result, the spin-glass-type regime appears from the critical state, as observed by the spin-glass correlators [39]. On the the hand for the mixed states, decoherence operation is nothing but measurements without monitoring outcomes, and it generates the spin-glass type symmetry, i.e, the SWSSB. Then, similarly to the ground state in the regime II of the qAT model, the glass property eliminates the ordinary LROs in the upper and lower chains, $\chi^{I}_{u}$ and $\chi^{I}_{Z,st}\sim 0$, but the glassy $Z_{2}$ order remains, which is captured by $\chi^{II}_{ZZ}$. This phenomenon is expected to appear rather generally in various systems under decoherence. In this sense, the similarity of the ground-state phase diagram of the qAT model and that of the above mixed state comes from rather specific choice of the $\lambda$-terms and multiple-decoherence protocol. This is a highlight of the present study.
@@ VIII Summary and conclusion
In this work, utility and specific character of the filtering formulation in the doubled Hilbert formalism to treat decoherence effects have been clarified. This filtering formulation and the physical observable that we proposed in this work have broad applications to various spin models, such as XXZ chain. In particular, how the gapless ground state of the XXZ model without LROs is affected by decoherence is interesting [60]. Since the doubled system that we consider corresponds to a ladder spin system, the accumulated knowledge of pure-state ground state phase diagrams in various such spin systems gives us insight into existence of interesting mixed states that we do not know yet.
@@ Acknowledgements
Data availability
The data and code that support the main findings of this study are available on Zenodo [61].
@@ Appendix B Canonical $Z_{2}$ correlator in the doubled Hilbert space
For the 1D TFIM under decoherence, the canonical $Z_{2}$ correlator is given by ${\rm Tr}[\rho_{D}Z_{i}Z_{j}]$, employed as an order parameter characterizing the ordinary LRO in statistical mechanics. As we explained in the main text, this observable detects SSB of both the strong and weak symmetries. It is useful to have an expression corresponding to this in the doubled Hilbert space formalism to analyze mixed states. By using the Choi isomorphism formula for density matrix $\rho$  [34, 23],